# VIST3A: Text-to-3D by Stitching a Multi-view Reconstruction Network to a Video Generator

### v1

@@ 4.1 Experimental Setups
Target 3D models. We target last-generation foundational 3D vision models that have been trained on large-scale datasets, have demonstrated generality and reliable performance across diverse domains, and require only images as input. For our experiments, we select three representative state-of-the-art models: (1) MVDUSt3R (Tang et al., 2025) predicts pointmaps and Gaussian splats, (2) VGGT (Wang et al., 2025a) predicts pointmaps, depth maps and camera poses, and (3) AnySplat (Jiang et al., 2025) predicts Gaussian splats and camera poses.
Target video generators. Our primary video model is Wan 2.1 T2V large (Wan et al., 2025), a state-of-the-art text-to-video generator. To demonstrate the generality of VIST3A across different architectures, we additionally use several other latent video models, including CogVideoX (Yang et al., 2024b), SVD (Blattmann et al., 2023), and HunyuanVideo (Kong et al., 2024).
@@ 4.2 Main Results: 3D Generation
Baselines. Important baselines for text-to-3DGS are SplatFlow (Go et al., 2025a), Director3D (Li et al., 2024), Prometheus3D (Yang et al., 2025c), and VideoRFSplat (Go et al., 2025b). Additionally, we include Matrix3D-omni (Yang et al., 2025d), to our knowledge, the only other model that unifies generation and reconstruction in latent space.
Evaluation protocol. We evaluate text-to-3DGS models on three benchmarks: T3bench (He et al., 2023) for object-centric generation, SceneBench (Yang et al., 2025c) for scene-level synthesis, and DPG-bench (Hu et al., 2024) to assess adherence to long, detailed prompts. On T3bench and SceneBench, we render images and compute Imaging Quality and Aesthetic Quality scores as defined by VBench (Huang et al., 2024) to assess visual fidelity, CLIP score (Hessel et al., 2021) for text-prompt alignment, and Alignment, Coherence, and Style scores according to Wang et al. (2025c) as comprehensive quality metrics. We prefer to avoid traditional no-reference metrics like NIQE (Mittal et al., 2012b) and BRISQUE (Mittal et al., 2012a) that have sometimes been used in the context of 3D generation, but lack a meaningful connection to the conditional generation task (e.g., they can be gambled by always returning the same sharp and colorful, high-scoring image, independent of the prompt). For DPG-bench, we follow the suggested protocol (Hu et al., 2024), but upgrade from the originally proposed language models to the more capable, UnifiedReward LLM (based on Qwen 7B). Text-to-pointmap models are evaluated qualitatively, as no established benchmarks or baselines exist.
@@ 4.3 Main Results: Model Stitching
Evaluation protocol. For 3DGS models, we evaluate novel-view synthesis on RealEstate10K (Zhou et al., 2018), with 8 source and 4 target images. For 3D reconstruction models, we follow Pi3 (Wang et al., 2025d) and assess pointmap quality on ETH3D (Schöps et al., 2017), and camera pose estimation on RealEstate10K and ScanNet (Dai et al., 2017). Specifically, Accuracy (Acc.), Completion (Comp.), and Normal Consistency (N.C.) are used for pointmap estimation, while camera pose estimation is evaluated with Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) at 5° and their AUC up to 30°.
@@
Mean Med. Mean Med. Mean Med.
MVDUSt3R 0.400 0.291 0.376 0.159 0.805 0.905 98.66 12.91 42.34 0.015 0.019 0.691
VGGT 0.263 0.188 0.197 0.120 0.855 0.961 99.51 15.75 50.06 0.015 0.015 0.500
Hunyuan+MVDUSt3R 0.405 0.288 0.399 0.166 0.802 0.887 98.36 12.40 41.97 0.016 0.019 0.668
SVD+MVDUSt3R 0.410 0.310 0.387 0.168 0.804 0.899 98.12 12.67 41.69 0.016 0.020 0.690
CogVid+MVDUSt3R 0.412 0.281 0.387 0.157 0.781 0.888 98.29 12.36 41.96 0.016 0.019 0.680
@@ 4.4 Ablations
Figure 5: MSE and pointmap quality on ETH3D vs. to stitching layer. Lower MSE in the stitching layer correlates with better 3D reconstruction.
Impact of the stitching index (Sec 3.1). We pick the best layer for stitching according to a fairly simple criterion, namely the one that best supports a linear transfer of the encoder latents. To analyze the impact of this design, we train stitched decoders for the combination (Wan+VGGT) while varying the stitching index. In Fig. 5, we see that layers with lower stitching residual indeed yield better pointmaps, supporting the MSE of the linear stitching layer as our selection criterion. Notably, early layers tend to exhibit lower MSEs. It appears that the latents are more compatible with low-level features that retain fine details.
@@ Appendix A Extended related works
A line of recent works follows a pipeline design, chaining together multiple modules and models. Typically, the first stage generates multi-view images from text or a single input image, followed by a separate reconstruction model that lifts these views into a 3D representation (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024b; Li et al., 2023; Park et al., 2025), with large models such as LRM (Hong et al., 2023) often used for this step. However, since the generative and reconstruction stages are trained and executed independently, errors accumulate across these parts (e.g., view inconsistency, texture flicker). Moreover, such pipeline schemes are less robust to latent-space perturbations than approaches where generation and reconstruction are performed jointly in the same latent space (see Section E).
A second category of methods (Liu et al., 2024; Yu et al., 2024b; Gao et al., 2024; Sun et al., 2024; Wang et al., 2024a) also generates multi-view images before lifting them into 3D, but replaces large pretrained reconstruction models with per-scene optimization of NeRFs (Mildenhall et al., 2021) or 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). While this strategy avoids reliance on pretrained decoders, it remains prone to error accumulation and requires costly per-scene optimization, making inference slow and computationally expensive.
@@
A third line of works introduces progressive expansion and refinement pipelines (Yu et al., 2024a; Ni et al., 2025; Chen et al., 2025b; Fridman et al., 2023; Feng et al., 2025; Yu et al., 2025). Some adopt iterative warping and inpainting strategies (Yu et al., 2024a; Ni et al., 2025; Fridman et al., 2023; Yu et al., 2025), while others leverage video generative models to unfold 3D scenes in a progressive manner (Chen et al., 2025b; Feng et al., 2025). Beyond these, additional works propose elaborate multi-stage pipelines that further increase complexity (Yang et al., 2025b; Ost et al., 2025). However, such designs are overly complex and suffer from slow inference.
Alignment for text-to-2D models.
Diffusion (Ham et al., 2025; Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2020) and flow-matching models (Liu et al., 2023) have achieved remarkable success in 2D generation tasks across both image and video domains, and also depth estimation (Ke et al., 2025). Building upon these advances, numerous studies have explored improvements in model architectures (Lee et al., 2024; Go et al., 2023b; Park et al., 2024; 2023), loss weighting strategies (Go et al., 2023a), and timestep or noise-level sampling schemes (Kim et al., 2024). Leveraging these developments, a variety of foundational 2D generative models for images (Rombach et al., 2022) and videos (Wan et al., 2025; Yang et al., 2024b) have recently emerged.
@@
Furthermore, recent studies have explored several strategies for aligning pretrained text-to-2D models with human preferences: (1) direct fine-tuning with scalar rewards (Clark et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025), (2) Reward Weighted Regression (RWR) (Peng et al., 2019; Lee et al., 2023), (3) Direct Preference Optimization (DPO) (Rafailov et al., 2023; Yang et al., 2024a), and (4) PPO-based policy gradients (Black et al., 2024; Fan et al., 2023; Liu et al., 2025). In this work, we adopt direct fine-tuning, which uses gradient-based feedback to align the generative model with the stitched decoder, ensuring that the resulting latents yield high-quality, 3D-consistent outputs.
Concurrent works on interchanging parts of VAEs.
@@
Concurrently,  Chen et al. (2025a) explores replacing pretrained VAE encoders with other pretrained visual encoders that extract semantic representations. However, their approach overlooks the compatibility between the VAE latent space and the representation space of the substituted vision encoder, which consequently requires extensive retraining to achieve alignment. In contrast, we explicitly measure the similarity between the VAE latent space and the representations of each layer in 3D models, and stitch the most linearly transferable layer into the latent space. As a result, the stitched model achieves seamless integration without requiring extensive retraining.
Appendix B Methodology Details and Its Implementation
In this section, we provide additional details about the methodology behind VIST3A, extending the description given in Section 3. We first elaborate on the architectural and training aspects of our stitching method in Section B.1, including the stitching layers and loss functions used for MVDUSt3R (Tang et al., 2025), VGGT (Wang et al., 2025a), and AnySplat (Jiang et al., 2025). Section B.2 then details the direct reward finetuning methodology, outlining the reward formulations and their implementation for each 3D model (VGGT, AnySplat, and MVDUSt3R).
@@ B.1 Model Stitching
Stitching layer.
We implement the stitching layer $\mathbf{S}$ as a single Conv3D layer. Relying only on Conv3D parameters to align the spatial and temporal dimensions between the latent and the features from $F_{k+1:l}$ can result in unnatural configurations, such as excessively large padding size. To address this, we first interpolate the latent representation to the target dimensions and then apply Conv3D, which provides a cleaner alignment of spatial and temporal dimensions. This design still admits a closed-form expression of the stitching layer, as shown in Eq. 2.
Loss function for each 3D model.
We train the stitched VAE using an $\ell_{1}$ loss between its outputs and those of the original 3D model. Since 3D model outputs often consist of multiple components, we compute the $\ell_{1}$ loss for each component separately and then aggregate them with a weighted sum. Assigning equal weights can destabilize training and even cause divergence, since some components (e.g., confidence terms) have much larger scales than others. To mitigate this, we reweight the component losses to approximately balance their scales. The specific weighting strategy is adapted to each 3D model as follows:
@@
Hyperparameters and implementation details.
For the stitching layer $\mathbf{S}$, we adopt a single 3D convolution with kernel size, stride, and padding chosen to align the latent features from the video VAE with the representation space of each 3D model:
@@ B.2 Direct Reward Finetuning
Reward details.
We combine CLIP-based scores and HPSv2.1 human preference scores to construct rewards for both multi-view image quality and 3D representation quality. Specifically, we use DFN (Fang et al., 2024) as the CLIP model and HPSv2.1 (Wu et al., 2023). Given an image $I$ and its associated prompt $c$, we denote the HPSv2.1 score as $s_{\text{hps}}$ and the DFN CLIP score as $s_{\text{clip}}$. The quality reward is then defined as
@@
Hyperparameter in sampling process.
For generating samples required in the $T_{1},T_{2}]$ direct reward tuning stage, we set $T_{1}=10$ and $T_{2}=50$ in Algorithm 1, ensuring that the number of diffusion steps is smaller than the typical steps in inference. The number of gradient-enabled steps is set to $K=2$ to reduce memory consumption during training. For scheduling, we adopt the default scheduler from Wan 2.1 (Wan et al., 2025).
@@ C.1 Training Setup
Setup for stitching layer search.
To identify the stitching layer, we rely on representations from the feedforward 3D model and the corresponding latents computed on the same dataset. Specifically, we utilize a subset of the DL3DV dataset, comprising 200 scenes for VGGT, 800 scenes for AnySplat, and 3,200 scenes for MVDUSt3R, with only 13 views per scene used for the search. We limit our search to the encoder layers of each model, as we observe that MSE values consistently increase within deeper layer indices.
Setup for stitched VAE finetuning.
We train on a combination of the DL3DV and ScanNet datasets, defining one epoch as a full pass over DL3DV and two passes over ScanNet. For each training iteration, a number of scenes are sampled according to the batch size. From each selected scene, we randomly sample 9 or 13 views to serve as input samples for training. The models are trained for 50 epochs in total. The batch sizes are set to 12 for VGGT, 24 for MVDUSt3R, and 12 for AnySplat. The learning rate is fixed at $2\times 10^{-4}$ for all models with cosine decay scheduling and 500-step warmup. For training, we use AdamW (Loshchilov & Hutter, 2017), apply gradient clipping with a norm threshold of 1.0, and use gradient checkpointing on each stitched VAE block to reduce memory consumption. In addition to LoRA parameters, for AnySplat and VGGT, we also finetune register tokens and class tokens. This is necessary because we remove the earlier layers that originally process these tokens into intermediate representations, requiring adaptation of the token handling mechanism. We further utilize gradient checkpointing for every stitched VAE block.
Setup for generative model finetuning.
We finetune the generative models using only the DL3DV dataset. For generative loss computation, we use a batch size of 12 with 13 views per scene. Reward calculation uses a prompt batch size of 4, with 13 views for AnySplat and MVDUSt3R, and 9 views for VGGT. We again adopt AdamW with a learning rate of $1\times 10^{-4}$, apply gradient clipping at a 0.1 norm, and train LoRA parameters with rank 8 and alpha 16. Gradient checkpointing is enabled for all model blocks to reduce memory usage.

### v2

@@ 4.1 Experimental Setups
Target 3D models. We target last-generation foundational 3D vision models that have been trained on large-scale datasets, have demonstrated generality and reliable performance across diverse domains, and require only images as input. For our experiments, we select three representative state-of-the-art models: (1) MVDUSt3R (Tang et al., 2025) predicts pointmaps and Gaussian splats, (2) VGGT (Wang et al., 2025b) predicts pointmaps, depth maps and camera poses, and (3) AnySplat (Jiang et al., 2025) predicts Gaussian splats and camera poses.
Target video generators. Our primary video model is Wan 2.1 T2V large (Wan et al., 2025), a state-of-the-art text-to-video generator. To demonstrate the generality of VIST3A across different architectures, we additionally use several other latent video models, including CogVideoX (Yang et al., 2025e), SVD (Blattmann et al., 2023), and HunyuanVideo (Kong et al., 2024).
@@ 4.2 Main Results: 3D Generation
Baselines. Important baselines for text-to-3DGS are SplatFlow (Go et al., 2025a), Director3D (Li et al., 2024b), Prometheus3D (Yang et al., 2025c), and VideoRFSplat (Go et al., 2025b). Additionally, we include Matrix3D-omni (Yang et al., 2025d), to our knowledge, the only other model that unifies generation and reconstruction in latent space.
Evaluation protocol. We evaluate text-to-3DGS models on three benchmarks: T3bench (He et al., 2023) for object-centric generation, SceneBench (Yang et al., 2025c) for scene-level synthesis, and DPG-bench (Hu et al., 2024) to assess adherence to long, detailed prompts. On T3bench and SceneBench, we render images and compute Imaging Quality and Aesthetic Quality scores as defined by VBench (Huang et al., 2024) to assess visual fidelity, CLIP score (Hessel et al., 2021) for text-prompt alignment, and Alignment, Coherence, and Style scores according to Wang et al. (2025d) as comprehensive quality metrics. We prefer to avoid traditional no-reference metrics like NIQE (Mittal et al., 2012b) and BRISQUE (Mittal et al., 2012a) that have sometimes been used in the context of 3D generation, but lack a meaningful connection to the conditional generation task (e.g., they can be gambled by always returning the same sharp and colorful, high-scoring image, independent of the prompt). For DPG-bench, we follow the suggested protocol (Hu et al., 2024), but upgrade from the originally proposed language models to the more capable, UnifiedReward LLM (based on Qwen 7B). Text-to-pointmap models are evaluated qualitatively, as no established benchmarks or baselines exist.
@@
Table 4: User study results. Participants rank five methods in terms of text alignment and visual quality of rendered videos from generated 3DGS (lower average rank is better).
Method Text Alignment (Avg. Rank $\downarrow$) Visual Quality (Avg. Rank $\downarrow$)
Director3D 3.03 2.99
SplatFlow 3.38 3.88
Prometheus3D 3.25 3.71
VideoRFSplat 2.74 2.92
[…]
Human evaluation. To further validate the effectiveness of VIST3A, we conduct a user study where we compare it against four other methods: Director3D, SplatFlow, Prometheus3D, and VideoRFSplat. A total of 28 participants evaluated 14 randomly selected samples drawn from T3Bench, SceneBench, and DPG-Bench, ranking each method according to two criteria: (1) Text Alignment and (2) Visual Quality of videos rendered from the generated 3DGS. As shown in Table 4, VIST3A achieves the best performance (lowest average rank) on both criteria. Notably, participants rank VIST3A as the top method in >68% of cases for text alignment and >87% for visual quality, underscoring its superiority in generating high-fidelity, semantically consistent 3D scenes.
@@ 4.3 Main Results: Model Stitching
Evaluation protocol. For 3DGS models, we evaluate novel-view synthesis on RealEstate10K (Zhou et al., 2018), with 8 source and 4 target images. For 3D reconstruction models, we follow Pi3 (Wang et al., 2026) and assess pointmap quality on 7Scenes (Shotton et al., 2013) and ETH3D (Schöps et al., 2017), and camera pose estimation on RealEstate10K and ScanNet (Dai et al., 2017). Specifically, Accuracy (Acc.), Completion (Comp.), and Normal Consistency (N.C.) are used for pointmap estimation, while camera pose estimation is evaluated with Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) at 5° and their AUC up to 30°.
@@
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
MVDUSt3R 0.026 0.011 0.030 0.013 0.730 0.838 0.400 0.291 0.376 0.159 0.805 0.905 98.66 12.91 42.34 0.015 0.019 0.691
VGGT 0.020 0.008 0.029 0.016 0.694 0.790 0.263 0.188 0.197 0.120 0.855 0.961 99.51 15.75 50.06 0.015 0.015 0.500
Hunyuan+MVDUSt3R 0.027 0.012 0.032 0.012 0.699 0.806 0.405 0.288 0.399 0.166 0.802 0.887 98.36 12.40 41.97 0.016 0.019 0.668
SVD+MVDUSt3R 0.026 0.011 0.030 0.013 0.727 0.834 0.410 0.310 0.387 0.168 0.804 0.899 98.12 12.67 41.69 0.016 0.020 0.690
CogVid+MVDUSt3R 0.028 0.012 0.033 0.014 0.699 0.808 0.412 0.281 0.387 0.157 0.781 0.888 98.29 12.36 41.96 0.016 0.019 0.680
@@ 4.4 Ablations
Figure 5: MSE and pointmap quality on 7-Scenes vs. to stitching layer. Lower MSE in the stitching layer correlates with better 3D reconstruction.
Effectiveness of MSE for finding stitching layer (Sec 3.1).
We pick the best layer for stitching according to a fairly simple criterion, namely the one that best supports a linear transfer of the encoder latents. To analyze the impact of this design, we train stitched decoders for the combination (Wan+VGGT) while varying the stitching index. In Fig. 5, we see that layers with lower stitching residual indeed yield better pointmaps, supporting the MSE of the linear stitching layer as our selection criterion.
This empirical trend is also consistent with existing theory: Theorem 1 in Insulla et al. (2025) shows that the stitching risk of the hybrid network, obtained by connecting the source model’s early layers $f_{1}$ with the target model’s latter layers $g_{2}$ via a linear map $S_{1,2}$, is upper-bounded by the MSE at the stitching layer,
$\mathbb{E}\big[\|g_{2}(S_{1,2}f_{1})(x)-g_{2}(f_{2})(x)\|^{2}\big]\leq\kappa_{2}^{2}\,\mathbb{E}\big[\|S_{1,2}f_{1}(x)-f_{2}(x)\|^{2}\big],$ (4)
where $\kappa_{2}$ is the Lipschitz constant of $g_{2}$. Thus, an MSE is related to the upper bound on the stitching error, supporting our use of MSE.
Figure 6: CKA visualization.
Furthermore, motivated by the observation of Insulla et al. (2025) that the right-hand side of Eq. 4 takes a similar form of kernel alignment, we investigate whether CKA (Kornblith et al., 2019) can track the trend of final performance. Figure 6 reports the CKA between the latent representation of the Wan VAE and the representations at different layer indices of VGGT (larger values indicate greater similarity). As shown, CKA captures the overall degradation in performance as the layer index increases; however, it is less precise than MSE in identifying the best layer, failing to capture that the best performance is achieved at layer 2. These results suggest that, in our setting, MSE is a more reliable indicator of transferability than CKA.
@@
Additional results in Appendix E demonstrate that VIST3A inherits prompt-based camera control from the video backbone (e.g., responding to instructions like ”aerial droneshot”), that the stitching analysis generalizes across multiple VAE architectures, and that our finetuning does not degrade — and even slightly improves — video generation quality as measured by VBench.
Additional results. In Appendix E, we further show that VIST3A inherits prompt-based camera control from the video backbone, and the stitching layer analysis generalizes across multiple VAE architectures, and our finetuning does not degrade video generation quality on VBench.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by an unrestricted gift from Google, and by a grant from the Swiss National Supercomputing Centre within the Swiss AI Initiative (project a144). We thank Shengyu Huang for insightful discussions and feedback on the paper, particularly regarding the relationship to Representation Autoencoders. We are also grateful to Yongwei Chen for help with reviewing, validating, and cleaning the code.
@@ Appendix A Extended related works
Pipeline-based 3D generation. A common design pattern chains multiple modules sequentially: a first stage generates multi-view images from text or a single image, and a second stage lifts those views into a 3D representation (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024b; Li et al., 2024a; Park et al., 2025). The reconstruction step is handled either by a large feedforward model such as LRM (Hong et al., 2024), or by per-scene optimization of NeRFs (Mildenhall et al., 2021) or 3DGS (Kerbl et al., 2023) (Liu et al., 2024; Yu et al., 2025b; Gao et al., 2024; Sun et al., 2025; Wang et al., 2025a). In both cases, the generative and reconstruction stages are trained independently, making the overall system prone to error accumulation (e.g., view inconsistency, texture flicker) and less robust to latent-space perturbations than joint formulations in latent space (See Section E). The per-scene optimization variant additionally incurs substantial inference cost.
@@
A separate line of work pursues progressive expansion and refinement (Yu et al., 2024; Ni et al., 2025; Chen et al., 2025; Fridman et al., 2023; Feng et al., 2025; Yu et al., 2025a). Some methods adopt iterative warping and inpainting (Yu et al., 2024; Ni et al., 2025; Fridman et al., 2023; Yu et al., 2025a), while others leverage video generative models to unfold 3D scenes progressively (Chen et al., 2025; Feng et al., 2025). Beyond these, additional works propose elaborate multi-stage pipelines that further increase complexity (Yang et al., 2025b; Ost et al., 2025). However, such designs are overly complex and suffer from slow inference.
@@
Alignment for text-to-2D models. Several strategies have been proposed to align pretrained text-to-image models with human preferences: direct finetuning with scalar rewards (Clark et al., 2024; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025), Reward Weighted Regression (Peng et al., 2019; Lee et al., 2023), Direct Preference Optimization (Rafailov et al., 2023; Yang et al., 2024), and PPO-based policy gradients (Black et al., 2024; Fan et al., 2023; Liu et al., 2025). Among these, direct reward finetuning offers gradient-based feedback that propagates through the full denoising trajectory, which makes it a natural fit for our setting, where the reward depends on the decoded 3D output.
@@
Concurrent works on interchanging parts of VAEs. Concurrently,  Chen et al. (2026) replace pretrained VAE encoders with visual encoders that extract semantic representations. Their approach does not explicitly account for the compatibility between the substituted encoder’s representation space and the downstream decoder, and consequently requires extensive retraining. In our framework, we measure the linear transferability between the VAE latent space and each layer of the 3D model, and stitch at the most compatible layer, which yields a well-initialized decoder that needs only lightweight finetuning.
Appendix B Methodology Details and Implementation
@@ B.1 Model Stitching
Stitching layer. We implement the stitching layer $\mathbf{S}$ as a single Conv3D layer. Relying only on Conv3D parameters to align the spatial and temporal dimensions between the latent and the features from $F_{k+1:l}$ can result in unnatural configurations, such as excessively large padding size. To address this, we first interpolate the latent representation to the target dimensions and then apply Conv3D, which provides a cleaner alignment of spatial and temporal dimensions. This design still admits a closed-form expression of the stitching layer, as shown in Eq. 2.
Loss function for each 3D model. We train the stitched VAE using an $\ell_{1}$ loss between its outputs and those of the original 3D model. Since 3D model outputs often consist of multiple components, we compute the $\ell_{1}$ loss for each component separately and then aggregate them with a weighted sum. Assigning equal weights can destabilize training and even cause divergence, since some components (e.g., confidence terms) have much larger scales than others. To mitigate this, we reweight the component losses to approximately balance their scales. The specific weighting strategy is adapted to each 3D model as follows:
@@
Hyperparameters and implementation details. For the stitching layer $\mathbf{S}$, we adopt a single 3D convolution with kernel size, stride, and padding chosen to align the latent features from the video VAE with the representation space of each 3D model:
@@ B.2 Direct Reward Finetuning
Reward details. We combine CLIP-based scores and HPSv2.1 human preference scores to construct rewards for both multi-view image quality and 3D representation quality. Specifically, we use DFN (Fang et al., 2024) as the CLIP model and HPSv2.1 (Wu et al., 2023). Given an image $I$ and its associated prompt $c$, we denote the HPSv2.1 score as $s_{\text{hps}}$ and the DFN CLIP score as $s_{\text{clip}}$. The quality reward is then defined as
@@
Hyperparameter in sampling process. For generating samples required in the $T_{1},T_{2}]$ direct reward tuning stage, we set $T_{1}=10$ and $T_{2}=50$ in Algorithm 1, ensuring that the number of diffusion steps is smaller than the typical steps in inference. The number of gradient-enabled steps is set to $K=2$ to reduce memory consumption during training. For scheduling, we adopt the default scheduler from Wan 2.1 (Wan et al., 2025).
@@ C.1 Training Setup
Setup for stitching layer search. To identify the stitching layer, we rely on representations from the feedforward 3D model and the corresponding latents computed on the same dataset. Specifically, we utilize a subset of the DL3DV dataset, comprising 200 scenes for VGGT, 800 scenes for AnySplat, and 3,200 scenes for MVDUSt3R, with only 13 views per scene used for the search. We limit our search to the encoder layers of each model, as we observe that MSE values consistently increase within deeper layer indices.
Setup for stitched VAE finetuning. We train on a combination of the DL3DV and ScanNet datasets, defining one epoch as a full pass over DL3DV and two passes over ScanNet. For each training iteration, a number of scenes are sampled according to the batch size. From each selected scene, we randomly sample 9 or 13 views to serve as input samples for training. The models are trained for 50 epochs in total. The batch sizes are set to 12 for VGGT, 24 for MVDUSt3R, and 12 for AnySplat. The learning rate is fixed at $2\times 10^{-4}$ for all models with cosine decay scheduling and 500-step warmup. For training, we use AdamW (Loshchilov and Hutter, 2019), apply gradient clipping with a norm threshold of 1.0, and use gradient checkpointing on each stitched VAE block to reduce memory consumption. In addition to LoRA parameters, for AnySplat and VGGT, we also finetune register tokens and class tokens. This is necessary because we remove the earlier layers that originally process these tokens into intermediate representations, requiring adaptation of the token handling mechanism. We further utilize gradient checkpointing for every stitched VAE block.
Setup for generative model finetuning. We finetune the generative models using only the DL3DV dataset. For generative loss computation, we use a batch size of 12 with 13 views per scene. Reward calculation uses a prompt batch size of 4, with 13 views for AnySplat and MVDUSt3R, and 9 views for VGGT. We again adopt AdamW with a learning rate of $1\times 10^{-4}$, apply gradient clipping at a 0.1 norm, and train LoRA parameters with rank 8 and alpha 16. Gradient checkpointing is enabled for all model blocks to reduce memory usage.